OpenStudy (crashonce):

If 1+2+3+...+r=n^2, where n is an integer less than 100, the possible values of r is/are?

jimthompson5910 (jim_thompson5910):

use the formula 1+2+3+...+r = r*(r+1)/2

jimthompson5910 (jim_thompson5910):

seems odd how they give you 2 variables, you sure it's not 1+2+3+...+n=n^2

OpenStudy (crashonce):

the options are 1, 8, or 49

jimthompson5910 (jim_thompson5910):

so see what you get when you compute 1+2+3+...+8 and also 1+2+3+...+49

jimthompson5910 (jim_thompson5910):

when r = 1, you have the trivial sum of just 1 (you don't have any other addends)

OpenStudy (crashonce):

336 and 2401

OpenStudy (crashonce):

yep

jimthompson5910 (jim_thompson5910):

how are you getting 336?

jimthompson5910 (jim_thompson5910):

and 2401

OpenStudy (crashonce):

sorry 36 lol (1+...+8) and 1225 is (1+...+49) sorry lol

jimthompson5910 (jim_thompson5910):

so the sum from 1 to 8 gives you a number that's a perfect square n^2 = 36 which means n = 6

OpenStudy (crashonce):

yep

OpenStudy (crashonce):

35^2=1225

jimthompson5910 (jim_thompson5910):

and summing from 1 to 49 gives n^2 = 1225 which means n = 35, yes

OpenStudy (crashonce):

yes so its 8 and 49

jimthompson5910 (jim_thompson5910):

so 1, 8, 49 look good

OpenStudy (crashonce):

could you help me with a few more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (crashonce):

ill post it